OpenStudy (mathmath333):

the sum of the second and the fifth term of an AP is 8 and that of the third and seventh is 14.find the eleventh term

OpenStudy (ikram002p):

\(a_2 +a_5=8\) \(a_3+a_7=14\) ok could you tell whats the \(a_n\) term in any AP ?

OpenStudy (mathmath333):

a(n)=a+(n-1)d

OpenStudy (ikram002p):

thus \(a_2=a_0+d\) \(a_5=a_0+4d\) \(a_0+d+a_0+4d=8\) \(2a_0+5d=8....eq (1)\) \(a_3=a_0+2d\) \(a_7=a_0+6d\) \(a_0+2d+a_0+6d=14\) \(2a_0+7d=14....eq (2)\) solve eq 1 and 2 for \(a_0\) and d

OpenStudy (mathmath333):